OpenStudy (anonymous):

You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly?

OpenStudy (anonymous):

P(correct) = 1/4 P(wrong) = 3/4 P(3 correct) = (5C3) (1/4)^3 (3/4)^2